OpenStudy (anonymous):

How can I rewrite (x^8/5 * x^4/5) / (x^2/5)?

OpenStudy (anonymous):

Its: \[\frac{ x ^{8/5} * x^{4/5} }{ x ^{2/5}}\]

OpenStudy (anonymous):

nevermind :3 its (x^(8/5) * x^(4/5)) / x^(2/5)

OpenStudy (anonymous):

\[\frac{8}{5}+\frac{4}{5}-\frac{2}{5}\] will be your exponent

OpenStudy (anonymous):

o.o What? @satellite73